# Bourbaki degree of pairs of projective surfaces

### v2

@@ 3.3 Locally free resolutions and the Bourbaki scheme
We may check computationally that the saturation of the ideal of $(2\times 2)$-minors coincides with the saturation of the annihilator ideal of $\mathcal{Q}_{\sigma}$ and with the saturation of the 0-th Fitting ideal, therefore over this subset $\mathcal{Q}_{\sigma}$ has rank one. The $1$-dimensional part of this scheme has $3$ irreducible components, two lines $L_{1}=V(x_{0},x_{3}),L_{2}=V(x_{2},x_{3})$ and a quadric plane curve $Q=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})$. Moreover, $L_{1}$ and $Q$ have multiplicity two structure, adding to $m(\sigma)=5$.
This can also be obtained from the minimal free resolution of ${\mathcal{T}}_{\sigma}$:
@@
so that $e=3$, $\operatorname{Bour}(\sigma)=4$ and $m(\sigma)=5$, with $c_{3}({\mathcal{T}}_{\sigma})=8$. This implies that $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$ but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$.
@@
considered in [11, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\xrightarrow{L}\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\xrightarrow{N}\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
So we compute $e=3$ and $m(\sigma)=7$. This may also be seen from the primary decomposition of $\operatorname{supp}(\mathcal{Q}_{\sigma})$, which is composed of three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$, where the last line has a multiplicity $5$ structure, and at every line the rank of $\mathcal{Q}_{\sigma}$ is one.
@@
Here, although the 0-th Fitting ideal is different from the annihilator ideal of $\mathcal{Q}_{\sigma}$, their one-dimensional component coincide with a multiplicity five structure along the line $V(x_{3},x_{0})$, and thus $m(\sigma)=5$.
@@
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as 16, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
@@ 4 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 28. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example [4]).
As explored in [11, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
@@
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
@@
    ([4, $2.3$, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
@@
Theorem B. Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
@@
[18, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree $2$ and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
@@
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results above to show that the only possible discrete invariant is $c_{3}({\mathcal{T}}_{\sigma})=8$, although the existence of such an example remains an open question.
Indeed, from [19, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 7 we obtain $c_{3}({\mathcal{T}}_{\sigma})=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
@@
Using the classification of $\mu$-semistable reflexive sheaves on $\mathbb{P}^{3}$ with small $c_{2}$, we are able to obtain a full classification of the cases of pencils of cubics with $m(\sigma)=7$ and $\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$ (see 42).
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$. If $m(\sigma)=3$, then $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})\geq 2$, and in particular ${\mathcal{T}}_{\sigma}$ must be $\mu$-stable.
@@
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Section 3.1, $(c)$, since $m(\sigma)=3\neq 7$, $\sigma$ is incompressible, and from [16, Theorem 1.1], we obtain $c_{3}\leq 4$ (for a more general formula, see 45). On the other hand, using Remark 7 we have $c_{3}({\mathcal{T}}_{\sigma})=6-2a$, so $a\in\{1,2,3\}$.
@@
which cannot coincide for the values $a\in\{1,2,3\}$, hence we obtain a contradiction. ∎

### v3

@@ 2.3 Free resolutions and the Bourbaki scheme
The scheme $(\Xi_{\sigma})_{1}$ has three irreducible components, given by two lines and a plane quadric curve:
$L_{1}$ $=V(x_{0},x_{3})\doteq V(\mathfrak{p}_{1})$
$L_{2}$ $=V(x_{2},x_{3})=V(\mathfrak{p}_{2})$
$Q$ $=V(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})=V(\mathfrak{p}_{3}),$
Now, we proceed to computing the length of the cokernel module $Q$ over each prime ideal $\mathfrak{p}_{i}$, in order to show that $m(\sigma)=5$ using the associativity formula.
Starting with the prime ideal $\mathfrak{p}_{1}=(x_{0},x_{3})$, we note that the element $3x_{2}^{2}\in R_{\mathfrak{p}_{1}}$ is invertible, and hence the sum $v=3x_{2}^{2}+x_{1}x_{3}$ is also invertible, as $x_{1}x_{3}\in\mathfrak{p}_{1}$. Let $u=v^{-1}$, so we perform two elementary operations on the matrix $\nabla\sigma$ over $\mathfrak{p}_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{1}}$ $\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&x_{2}x_{3}&v&x_{1}x_{2}+3x_{3}^{2}\end{pmatrix}$
$\sim\begin{pmatrix}2x_{0}x_{2}+x_{1}x_{3}&x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3}&0&x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2})\\ 0&ux_{2}x_{3}&1&u(x_{1}x_{2}+3x_{3}^{2})\end{pmatrix}$.
The first transform is to add the scaled second row (by $ux_{0}^{2}$) to the first, and then we scale the second row by $u$. Then, the third basis vector goes via $\nabla\sigma$ to $(0,1)\in R_{\mathfrak{p}_{1}}^{2}$, and therefore we may compute the cokernel from the quotient
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}$ $\simeq\frac{R_{\mathfrak{p}_{1}}}{(2x_{0}x_{2}+x_{1}x_{3},x_{0}x_{3}-ux_{0}^{2}x_{2}x_{3},x_{0}x_{1}+3x_{3}^{2}-ux_{0}^{2}(x_{1}x_{2}+3x_{3}^{2}))}\simeq\frac{R_{\mathfrak{p}_{1}}}{I},$
where $I=(f_{1},f_{2},f_{3})$ are the generators above. We may factor $f_{2}=x_{0}x_{3}(1-ux_{0}x_{2})$, so that since $(1-ux_{0}x_{2})$ is a unit in the local ring, we conclude $x_{0}x_{3}\in I$. From this, multiplying $x_{3}f_{1}$ we conclude that $x_{1}x_{3}^{2}\in I$, and since $x_{1}$ is also a unit, we obtain $x_{3}^{2}\in I$, and analogously taking $x_{0}f_{1}$ we obtain $x_{0}^{2}x_{2}\in I$, and since $x_{2}$ is a unit we get $x_{0}^{2}\in I$. Thus
$(x_{0}^{2},x_{0}x_{3},x_{3}^{2})\subset I.$
Now, from this, looking at $f_{3}$:
$f_{3}=x_{0}x_{1}+3x_{3}^{2}-x_{0}^{2}u(x_{1}x_{2}+3x_{3}^{2})\in I\Rightarrow x_{0}x_{1}\in I\Rightarrow x_{0}\in I,$
and looking at $f_{1}$, we obtain
$f_{1}=2x_{0}x_{2}+x_{1}x_{3}\in I\Rightarrow x_{1}x_{3}\in I\Rightarrow x_{3}\in I,$
so that $I=(x_{0},x_{3})$ and thus $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}}\simeq\kappa$, with $\operatorname{{length}}((\mathcal{Q}_{\sigma})_{\mathfrak{p}_{1}})=1$.
For the case $\mathfrak{p}_{2}=(x_{2},x_{3})$, let $u=(x_{0}^{2})^{-1}$, $q=(3x_{2}^{2}+x_{1}x_{3})$ and consider the elementary operations given by $R_{2}^{\prime}=R_{2}-uqR_{1}$ and then $R_{1}^{\prime}=uR_{1}$:
$(\nabla\sigma)_{\mathfrak{p}_{2}}$ $\sim\begin{pmatrix}u(2x_{0}x_{2}+x_{1}x_{3})&ux_{0}x_{3}&1&u(x_{0}x_{1}+3x_{3}^{2})\\ -uq(2x_{0}x_{2}+x_{1}x_{3})&x_{2}x_{3}-uqx_{0}x_{3}&0&x_{1}x_{2}+3x_{3}^{2}-uq(3x_{3}^{2}+x_{0}x_{1})\end{pmatrix}$.
The Jacobian matrix sends the third basis vector to $(1,0)\in R_{\mathfrak{p}_{2}}^{2}$, and thus we may compute:
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}$ $\simeq\frac{R_{\mathfrak{p}_{2}}}{(f_{1},f_{2},f_{3})}\doteq\frac{R_{\mathfrak{p}_{2}}}{I}$
where
$f_{1}$ $=(3x_{2}^{2}+x_{1}x_{3})(2x_{0}x_{2}+x_{1}x_{3})$
$f_{2}$ $=x_{3}(x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3})$
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3ux_{3}^{2}(3x_{2}^{2}+x_{1}x_{3})-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})$
First, we consider
$f_{2}-\frac{x_{3}}{x_{1}}f_{3}=\frac{3x_{3}^{3}}{x_{1}}(1-u(3x_{2}^{2}+x_{1}x_{3}))\in I,$
and thus $x_{3}^{3}\in I$. Then
$x_{3}f_{2}=x_{3}^{2}x_{2}(1-3ux_{2})-ux_{0}x_{1}x_{3}^{3}\in I\Rightarrow x_{3}^{2}x_{2}\in I,$
and moreover
$x_{3}f_{1}$ $=x_{2}x_{3}(6x_{0}x_{2}^{2}+3x_{1}x_{2}x_{3}+2x_{0}x_{1}x_{3})+x_{1}^{2}x_{3}^{3}\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}+x_{3}^{2}x_{2}(3x_{1}x_{2}+2x_{0}x_{1})\in I$
$\Rightarrow 6x_{0}x_{2}^{3}x_{3}\in I$
$\Rightarrow x_{2}^{3}x_{3}\in I.$
On the other hand,
$x_{2}x_{3}f_{3}$ $=-3ux_{3}^{3}q+3x_{2}x_{3}^{3}+x_{2}x_{3}(x_{1}x_{2}-ux_{0}x_{1}q)\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}-3ux_{0}x_{1}x_{2}^{3}x_{3}-ux_{0}x_{1}^{2}x_{2}x_{3}^{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}x_{3}\in I\Rightarrow x_{2}^{2}x_{3}\in I.$
Now, consider
$x_{3}f_{3}$ $=x_{1}x_{2}x_{3}+3x_{3}^{3}-uqx_{3}^{3}-uqx_{0}x_{1}x_{3}\in I$
$\Rightarrow x_{3}x_{2}(x_{2}-3ux_{0}x_{2}^{2}-ux_{0}x_{1}x_{3})\in I$
$\Rightarrow x_{1}x_{3}(x_{2}-ux_{0}x_{1}x_{3})\doteq A\in I,$
and
$x_{3}A\in I\Rightarrow x_{1}x_{3}^{2}x_{2}-ux_{0}x_{1}^{2}x_{3}^{2}\in I\Rightarrow x_{3}^{2}\in I.$
This simplifies
$f_{2}=x_{3}(x_{2}-uqx_{0})\in I\Rightarrow x_{2}x_{3}(1-3ux_{2})\in I\Rightarrow x_{2}x_{3}\in I.$
Considering
$f_{1}$ $=6x_{0}x_{2}^{3}+x_{2}x_{3}(3x_{1}x_{2}+2x_{0}x_{1})+x_{1}^{2}x_{3}^{2}\in I$
$\Rightarrow x_{2}^{3}\in I,$
and then
$x_{2}f_{3}$ $=x_{1}x_{2}^{2}+x_{2}(3x_{3}^{2}-3ux_{3}^{2}q)-uqx_{0}x_{1}x_{2}\in I$
$\Rightarrow x_{1}x_{2}^{2}-3ux_{0}x_{1}x_{2}^{3}-ux_{0}x_{1}^{2}x_{2}x_{3}\in I$
$\Rightarrow x_{1}x_{2}^{2}\in I\Rightarrow x_{2}^{2}\in I,$
simplifying
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-3uqx_{3}^{2}-ux_{0}x_{1}(3x_{2}^{2}+x_{1}x_{3})\in I$
$\Rightarrow x_{1}(x_{2}-ux_{0}x_{1}^{2}x_{3})\in I\Rightarrow x_{2}-ux_{0}x_{1}x_{3}\in I,$
so that $(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}\simeq\kappa\langle 1,x_{2}\rangle$, meaning $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{2}}=2$, since the equations $f_{1},f_{2}$ are trivialized after $x_{3}^{2},x_{2}^{2},x_{2}x_{3}\in I$.
Let us consider the case $\mathfrak{p}_{3}=(x_{0}-x_{2},2x_{2}^{2}+x_{1}x_{3})\doteq(g_{1},g_{2})$. One may be tempted to choose $u=(2x_{0}x_{2}+x_{1}x_{3})^{-1}$ and do the elementary operations $R_{1}^{\prime}=uR_{1}$, but
$2x_{0}x_{2}+x_{1}x_{3}=2x_{0}g_{1}+g_{2}.$
Take $u\doteq(x_{0}^{2})^{-1}\in R_{\mathfrak{p}_{3}}$. If $x_{0}^{2}\in\mathfrak{p}_{3}$, then:
$x_{0}\in\mathfrak{p}_{3}\Rightarrow x_{2}\in\mathfrak{p}_{3}\Rightarrow x_{1}x_{3}\in\mathfrak{p}_{3}$
and $\text{height}(\mathfrak{p}_{3})\geq 3$ would give a contradiction. Thus, we may perform the same elementary operations as we did for $\mathfrak{p}_{2}$, and thus
$(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{3}}\simeq\frac{R_{\mathfrak{p}_{3}}}{I}$
with generators $I=(f_{1},f_{2},f_{3})$ given as before. By the same reasoning, $(3x_{2}^{2}+x_{1}x_{3})\not\in\mathfrak{p}_{3}$, and thus
$f_{1}\in I\Rightarrow 2x_{0}x_{2}+x_{1}x_{3}\in I.$
From the previous description of the element above, it suffices to show $g_{i}\in I$ for some $i\in\{1,2\}$ to conclude $I=\mathfrak{p}_{3}R_{\mathfrak{p}_{3}}$. Let $A\doteq(g_{2}+x_{2}^{2})$, so that $u=A^{-1}$. We can rewrite $f_{3}$ as:
$f_{3}$ $=x_{1}x_{2}+3x_{3}^{2}-u(g_{2}+x_{2}^{2})(3x_{3}^{2}+x_{0}x_{1})$
$=3x_{3}^{2}(1-u(g_{2}+x_{2}^{2}))+x_{1}(x_{2}-ux_{0}(g_{2}+x_{2}^{2}))$
$=3x_{3}^{2}(1-1)+x_{1}(x_{2}-ux_{0})$
$=-x_{1}g_{1}\in I,$
and since $x_{1}\notin\mathfrak{p}_{4}$, $g_{1}\in I$ and $I=\mathfrak{p}_{4}R_{\mathfrak{p}_{4}}$, concluding $\operatorname{{length}}(\mathcal{Q}_{\sigma})_{\mathfrak{p}_{4}}=1$. From the associativity formula, we obtain
$m(\sigma)=1\cdot 1+2\cdot 1+1\cdot 2=5.$
Since $m(\sigma)\leq 6$, we conclude using the bounds from the next section that $e\geq 3$ The minimal free resolution of ${\mathcal{T}}_{\sigma}$ given using Macaulay2 is
@@
so that $e=3$ and $\operatorname{Bour}(\sigma)=4$. This means $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$, but we need $4$ syzygies to generate ${\mathcal{T}}_{\sigma}$ or, equivalently, the degree four Bourbaki scheme is not a complete intersection.
@@
considered in [undefj, Theorem 8.1]. From their proof, we know that ${\mathcal{T}}_{\sigma}(2)$ is a null correlation bundle. Therefore, we obtain $\operatorname{Bour}(\sigma)=2$, and a free resolution for ${\mathcal{T}}_{\sigma}$ is given by:
$0\rightarrow\mathcal{O}_{\mathbb{P}^{3}}(-5)\to\mathcal{O}_{\mathbb{P}^{3}}(-4)^{\oplus 4}\to\mathcal{O}_{\mathbb{P}^{3}}(-3)^{\oplus 5}\rightarrow{\mathcal{T}}_{\sigma}\rightarrow 0.$
So $e=3$ and $m(\sigma)=7$. The irreducible components of $(\Xi_{\sigma})_{1}$ are three lines $V(x_{2},x_{3}),V(x_{1}-x_{2},x_{3})$ and $V(x_{1},x_{2})$. Localizing the cokernel at the first two lines gives length one, so over last line should have length $5$.
@@
Here, the unique irreducible component of $(\Xi_{\sigma})_{1}$ is the line $V(x_{0},x_{3})$.
@@
This is an example with the same discrete invariants $(\operatorname{indeg}({\mathcal{T}}_{\sigma}),\operatorname{Bour}(\sigma),m(\sigma))$ and the same total Chern class as Example 15, and it is not $3$-syzygy (neither $\operatorname{gpdim}({\mathcal{T}}_{\sigma})=1$) as the previous case.
@@ 3 Extreme cases of low initial degree
In this section, we observe that a non-zero section $\nu\in H^{0}({\mathcal{T}}_{\sigma}(e))$ induces a sub-foliation by curves of degree $e+1$ of the foliation ${\mathcal{T}}_{\sigma}(1)$. We derive numerical restrictions for this behavior when the initial degree is extremely low $e\in\{1,2\}$ using the classification of foliations by curves in $\mathbb{P}^{3}$ of degrees one and two. From this main result, we conclude that ${\mathcal{T}}_{\sigma}$ is $\mu$-stable when $\sigma=(f,g)$ is a sequence with $d_{f}=1,d_{g}=2$ and $m(\sigma)=3$, see 27. We will review some of the theory of foliations by curves in $\mathbb{P}^{3}$ (see, for example, [undefc]).
As explored in [undefj, Section 9], the sheaf ${\mathcal{T}}_{\sigma}(1)$ defines a foliation with a corresponding short exact sequence:
@@
where $\Gamma_{\sigma}\subset\mathbb{P}^{3}$ is the singular scheme of the associated foliation. Then, assuming $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})$, there is a non-zero section of $\nu\in\text{Hom}(\mathcal{O}_{\mathbb{P}^{3}}(1-e),{\mathcal{T}}_{\sigma}(1))$, inducing the commutative diagram with exact rows below
@@
    ([undefk, Theorem $4$]) If $e=1$, then $W$ is either a 0-dimensional scheme of length $4$, a union of a line with a zero-dimensional scheme of length two or double line of genus 1-1. In either case, $\deg(W)\leq 2$ or $W$ is zero-dimensional.
@@
For completeness, we include an argument for $(b)$, from V. Cordeiro: using the short exact sequence
$0\to G^{\vee}\to\Omega^{1}_{\mathbb{P}^{3}}\to\mathcal{I}_{W}(1)\to 0,$
one obtains $4\leq c_{2}(G^{\vee})=11-\deg(W)\leq 11$ (see, for example, [undefc, 4.1, 4.2]). Now, if $G^{\vee}$ is stable, Bogomolov’s inequality says $c_{2}(G^{\vee})\geq 7$. Now, assuming that $G^{\vee}$ is not stable, $h^{0}(G^{\vee}(2))\neq 0$ and one may choose a section to form a sequence
$0\to\mathcal{O}_{\mathbb{P}^{3}}(-2)\to G^{\vee}\to\mathcal{I}_{Y}(-3)\to 0$
where $Y$ is the vanishing locus of this section. Hence, from additivity one gets $0\leq\deg(Y)=c_{2}(\mathcal{I}_{Y})=c_{2}(G^{\vee})-6$. From the first inequality, we conclude $c_{2}(G^{\vee})\geq 6$, and thus $\deg(W)\leq 5$.
Theorem B.
Let $\sigma=(f,g)$ be a normal sequence of polynomials of degrees $d_{f}+1,d_{g}+1$. Then:
@@
[undeft, 1.4-1.6] Let $B\subset\mathbb{P}^{3}$ be a curve of degree 22 and genus $p_{a}(B)=-1-a$, for $a\in\mathbb{Z}$. Then:
@@
The results above can be used to restrict the possible values of numerical invariants. For example, for normal pencils of cubics ($d_{f}=d_{g}=2$), $m(\sigma)=7$ and $e=\operatorname{indeg}({\mathcal{T}}_{\sigma})=1$, we obtain $\operatorname{Bour}(\sigma)=2$ and we can use the results to show the only possible third Chern class is $c_{3}({\mathcal{T}}_{\sigma})=8$.
Indeed, from [undefu, Theorem 3.8] for the reflexive sheaf ${\mathcal{T}}_{\sigma}(2)$ (unstable of order $r=1$), $c_{3}({\mathcal{T}}_{\sigma})\leq 10$, and using Remark 6 we obtain $c_{3}({\mathcal{T}}_{\sigma})=2(p_{a}(B)+5)=8-2a$, where $p_{a}(B)=-1-a$. From the short exact sequence
@@
In the next section, we show the above cannot happen either (see 43).
@@
Let $\sigma=(f,g)$ be a normal sequence of homogeneous polynomials such that $d_{f}=1$, $d_{g}=2$ and $m(\sigma)=3$. If ${\mathcal{T}}_{\sigma}$ is unstable, then the possible discrete invariants are $e=1$ and $c_{3}({\mathcal{T}}_{\sigma})=6$.
@@
Let us assume that $e=1$, so that $\operatorname{Bour}(\sigma)=2$. From Theorem A Theorem A, $\sigma$ is incompressible. Using Remark 6 we have for $\deg(B)=2$, $p_{a}(B)=-1-a$ yields $c_{3}({\mathcal{T}}_{\sigma})=6-2a\geq 0$, so $a\in\{-1,0,1,2,3\}$, but from
@@
which can only coincide for $a=0$. ∎
We note that we have no explicit construction for the unstable case above to occur as an example.